OpenStudy (anonymous):

Solve for X. 5/6 * X = 30 *

OpenStudy (anonymous):

Solve for X. 5/6 * X = 30 *

OpenStudy (anonymous):

36

OpenStudy (michele_laino):

here we have to multiply both sides by 6, so we get: \[6 \cdot \frac{5}{6}x = 30 \cdot 6\]

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

use this website, super easy http://www.geteasysolution.com/5/6*x=30

OpenStudy (michele_laino):

now, we have: \[6 \cdot \frac{5}{6} = \frac{6}{6} \cdot 5 = ...?\]

OpenStudy (anonymous):

5 6?

OpenStudy (michele_laino):

what is \(\Large \frac{6}{6}\) ?

OpenStudy (anonymous):

huan 1

OpenStudy (michele_laino):

ok! So, we can write: \[6 \cdot \frac{5}{6} = \frac{6}{6} \cdot 5 = 1 \cdot 5 = ...?\]

OpenStudy (anonymous):

5

OpenStudy (michele_laino):

correct! so, the left side, is: \[6 \cdot \frac{5}{6}x = 5x\] now, what is \(30 \cdot 6=...\) ?

OpenStudy (anonymous):

ok hmm 180

OpenStudy (michele_laino):

that's right! So, I can rewrite the original equation like below: \(5x=180\)

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

1 more

OpenStudy (michele_laino):

Please wait. now, I divide both sides by 5, so I get: \[\frac{5}{5}x = \frac{{180}}{5}\]

OpenStudy (anonymous):

oops

OpenStudy (michele_laino):

what is \[\frac{5}{5}=...?\]

OpenStudy (anonymous):

1

OpenStudy (michele_laino):

ok! So left side become: \[\frac{5}{5} \cdot x = 1 \cdot x = x\]

OpenStudy (michele_laino):

and, what is \(180/5=...?\)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

36

OpenStudy (anonymous):

:D

OpenStudy (michele_laino):

please use windows calculator :)

OpenStudy (anonymous):

36

OpenStudy (michele_laino):

that's right! So, we the requested answer, is: \(x=36\)

OpenStudy (anonymous):

ok next one :D

OpenStudy (michele_laino):

ok!